OpenStudy (anonymous):

If you were asked to convert 4.2 X 10^22 atoms of aluminum to moles, which of the following should you use for the conversion?

imqwerty (imqwerty):

\[\frac{ 4.2 \times 10^{22}}{ N_{A} }\] N_{A} is the avagadro number

imqwerty (imqwerty):

yea sure :)

imqwerty (imqwerty):

1mole of any substance has 6.022 x 10^23 particles of it

imqwerty (imqwerty):

so 1 mole of Al has 6.022 x 10^23 atoms

OpenStudy (mertsj):

well which one of those is true? is 6.023x10^23 atoms of Al equal to 13 moles Al

imqwerty (imqwerty):

now if 1 mole has 6.022 x 10^23 atom then how many mole will have 4.2 x 10^22 atoms :)

imqwerty (imqwerty):

can u calculate how many moles will have 1 atom?

imqwerty (imqwerty):

well just because it has that bad number 26.891... we can't say its wrng we have to give proper reasoning ok we know that 1mole has 6.022 x 10^23 particles we need to find how many moles will have 1 particle so 1st answer this- 6 bags have 3 chocolates how many bags will have 1 chocolate?

OpenStudy (wolf1728):

What's with the mass of aluminum? The origianl question just want number of moles.

imqwerty (imqwerty):

ur saying B is correct cause it has got mass/ nvm its ok :)

OpenStudy (wolf1728):

But the original question was just "how many moles is 4.2 x 10^22 atoms".

OpenStudy (wolf1728):

There are 6.023 x 10^23 atoms per mole. You have 4.2 x 10^22 atoms. 4.2 x 10^22 * 1 mole/6.023 x 10^23 atoms

imqwerty (imqwerty):

yes

OpenStudy (wolf1728):

Or number 4 in your list

OpenStudy (wolf1728):

sure

OpenStudy (wolf1728):

2.03 X 10^24 atoms * 1 mole/6.023 x 10^23

OpenStudy (wolf1728):

Look carefully at my answer. You don't multiply both numbers.

imqwerty (imqwerty):

number of moles = no of atoms /avagadro number :)

OpenStudy (wolf1728):

I made a graphic which shows the calculation more clearly:

imqwerty (imqwerty):

yes :D

OpenStudy (wolf1728):

B it is!!

OpenStudy (wolf1728):

okay 1 more question.

imqwerty (imqwerty):

np B)

imqwerty (imqwerty):

moles=mass/molar mass molar mass of silver=108

imqwerty (imqwerty):

yea correct :D

OpenStudy (wolf1728):

Remember to learn how to set it up correctly. What if you set it up as 107.9/68.3?

OpenStudy (wolf1728):

Yes you set it up the right way 68.3/107.9 but you will have to see why.

OpenStudy (wolf1728):

The important thing is if you know the correct way to set up the answer because I'm assuming you will be tested on this, right?

OpenStudy (wolf1728):

So then you will definitely have to learn how to set up the answer.

OpenStudy (wolf1728):

Well imqwerty and myself looked at the answers and we're sure they are correct.

OpenStudy (wolf1728):

Atomic mass of copper = 63.55 meaning there are 6.023 x 10^23 atoms for every 63.55 grams of copper. We need to know atoms in 68.7 grams of copper: 68.7 * 6.023 x 10^23 atoms / 63.55 grams copper = 6.511 x 10^23 atoms which is C which is your answer

OpenStudy (wolf1728):

Moles in 125 g of NI atomic mass Ni = 58.71 grams per mole 125 grams * 1 mole / 58.71 grams = 2.129 moles B (which is your answer)

imqwerty (imqwerty):

#( 1mole of silver=108gm=6.022x10^23 particles \[\frac{ 108\times 5.8 \times 10^{23} }{ 6.022 \times 10^{23}}=mass of (5.8 \times 10^{23})particles\]

imqwerty (imqwerty):

^#9

OpenStudy (wolf1728):

I'm working on number 8

OpenStudy (wolf1728):

It seems to me you just want correct answers as opposed to HOW to set up the answers. Question 8 - your answer is wrong. What is the mass of 25 atoms of mercury? (your answer)

imqwerty (imqwerty):

yes sofia :) thas correct

OpenStudy (wolf1728):

So what is your answer for #8?

imqwerty (imqwerty):

np :) sofia can u tell how did u find out the mass

imqwerty (imqwerty):

nvm :) tell ur method

OpenStudy (wolf1728):

Yes it is either B or C for #8

imqwerty (imqwerty):

yes

imqwerty (imqwerty):

ok :)

imqwerty (imqwerty):

yes :D

imqwerty (imqwerty):

np (:

OpenStudy (wolf1728):

Did you figure out #8 yet?

OpenStudy (wolf1728):

Yes it is C - very good!